Eclipse Solar Company operates two factories. The company applies factory overhead to jobs on the basis of machine hours in Factory 1 and on the basis of direct labor hours in Factory 2. Estimated factory overhead costs, direct labor hours, and machine hours are as follows:

Factory 1 Factory 2
Estimated factory overhead cost for fiscal
year beginning August 1 $468,920 $633,600
Estimated direct labor hours for year 8,800
Estimated machine hours for year 24,680
Actual factory overhead costs for August $37,590 $54,600
Actual direct labor hours for August 790
Actual machine hours for August 1,930

a. Determine the factory overhead rate for Factory 1.
$ per machine hour

b. Determine the factory overhead rate for Factory 2.
$ per direct labor hour

c. Journalize the entries to apply factory overhead to production in each factory for August. If an amount box does not require an entry, leave it blank.

Factory 1
Factory 2

d. Determine the balances of the factory overhead accounts for each factory as of August 31, and indicate whether the amounts represent overapplied factory overhead or underapplied factory overhead.

Factory 1 $
Factory 2 $
Answer #1

a. Determine the factory overhead rate for Factory 1.
$468920/24680 = 19 per machine hour

b. Determine the factory overhead rate for Factory 2.
$633600/8800 = 72 per direct labor hour

c. Journalize the entries to apply factory overhead to production in each factory for August. If an amount box does not require an entry, leave it blank.

Factory 1 Work in process (1930*19) 36670
Factory overhead 36670
Factory 2 Work in process (790*72) 56880
Factory overhead 56880

d. Determine the balances of the factory overhead accounts for each factory as of August 31, and indicate whether the amounts represent overapplied factory overhead or underapplied factory overhead.

Factory 1 $36670-37590 = 920 Under applied
Factory 2 56880-54600 = 2280 Over applied